How many grams of O2 will be formed from 3.76 grams of KCIO3


1
Expert's answer
2021-10-07T01:25:03-0400

2KClO3 ==> 2KCl + 3O2

Molar mass of KClO3 = 122.55g/mol

3.76/122.55= 0.031mol

Mole ratio KClO3:O2 = 2:3

= 0.031×3/2

= 0.046m

= 0.046× 31.998

= 1.47g